### v1

# Towards Fast Inference: Exploring and Improving Blockwise Parallel Drafts

Abstract
Despite the remarkable strides made by autoregressive language models, their potential is often hampered by the slow inference speeds inherent in sequential token generation. Blockwise parallel decoding (BPD) was proposed by Stern et al. (2018) as a way to improve inference speed of language models. In this paper, we make two contributions to understanding and improving BPD drafts. We first offer an analysis of the token distributions produced by the BPD prediction heads. Secondly, we use this analysis to inform algorithms to improve BPD inference speed by refining the BPD drafts using small n-gram or neural language models. We empirically show that these refined BPD drafts yield a higher average verified prefix length across tasks.
Language Model
1 Introduction
The landscape of natural language processing has been profoundly reshaped by recent advances in autoregressive language models (Brown et al., 2020; Wei et al., 2022; Radford et al., 2019; Raffel et al., 2020; Team et al., 2023). These models have shown remarkable proficiency across a range of text generation tasks, including applications like question answering (Rajpurkar et al., 2016) and summarization (Hermann et al., 2015). However, a significant obstacle to their wider application is high inference latency, particularly for extremely deep models with hundreds of billions of parameters (Hoffmann et al., 2022; Rae et al., 2021; Chowdhery et al., 2022). This latency, intrinsic to decoding with autoregressive language models (LMs), imposes considerable computational burdens and limits real-time deployment.
In response to these latency challenges, the field has seen a shift towards decoding methods aimed at reducing the inference latency in large language models (LLM). One promising development is the concept of blockwise parallel decoding (BPD) (Stern et al., 2018; Monea et al., 2023; Cai et al., 2024). Unlike autoregressive decoding, which generates one token at a time, BPD LMs are outfitted with a set of prediction heads, which propose and verify a draft, a block of subsequent tokens, in parallel. While BPD offers one solution to accelerated text generation, it also poses a challenge in ensuring that the proposed drafts are fluent and natural.
Figure 1: Illustration of two tokens that are decoded by autoregressive decoding vs. two tokens drafted by BPD.
BPD inference speed is a function of the time required to produce a BPD draft, and how well that draft matches the decode produced by the base LM (Figure 1). While standard autoregressive LMs generate tokens sequentially, ensuring that each generated token is consistent with all of the preceding tokens (e.g., ‘Messi’ is consistent with ‘Lionel’), BPD adopts a different strategy. At each time step, BPD LMs predict a draft of multiple tokens concurrently (‘Lionel’ and ‘Ronaldo’), each of which is predicted independently of the others. The inherent challenge in BPD drafts lies in maintaining consistency between tokens at each position during parallel generation. A good BPD drafter should prioritize congruous sequences, like ‘Lionel Messi’, over incongruous sequences, like ‘Lionel Ronaldo’, which are unlikely to be decoded by a strong LM. The focus of this work is improving the quality of BPD drafts, without adding additional model parameters that would need to be trained.
2 Our Contributions
In this paper, we first investigate properties made by the prediction heads of BPD LMs across several tasks; given these observations, we propose rescoring algorithms to produce higher quality BPD drafts.
2.1 Observations on BPD Drafts
Consecutive repetitions  All heads within a block make predictions independently in a BPD LM. Unsurprisingly, we observe that this leads to BPD drafts with significant token repetition across heads. Consecutive repetition is pervasive across tasks, ranging from 20% to 75% of all neighboring draft tokens, depending on the task (subsection 6.1).
Confidence of different heads  We analyze the distribution of probabilities within each softmax head. Our empirical analysis reveals a key property of BPD: the BPD drafter tends to be more confident with initial tokens, and becomes progressively less confident for subsequent tokens. We find that the confidence of BPD heads correlates strongly with the quality of the BPD drafter (subsection 6.2).
Oracle top-$k$ block efficiency  In the BPD algorithm (Algorithm 1), the most likely token at each head is generated as the draft. As highlighted before, this approach is prone to two issues: (1) there might be consecutive repetitions and (2) the model might not be confident about the prediction at some of the heads. We use block efficiency, the average number of draft tokens accepted during decoding, to measure the quality of a given drafter (Leviathan et al., 2023; Sun et al., 2023b). We ask if the block efficiency can be improved by considering the top-$k$ most likely tokens at each head. To measure the potential benefit of considering top-$k$ tokens, we measure the block efficiency of the oracle path through this top-$k$ lattice, oracle top-$k$ block efficiency, and show that there is significant headroom for improvement across tasks (subsection 6.3).
2.2 New Algorithms
Based on these observations, we propose two algorithms to leverage the top-$k$ predictions at each head and improve BPD latency. Neither of these algorithms require changes to the underlying BPD LM.
Local rescoring via neural model  Given the top-$k$ predictions at each head, we refine the BPD draft by using a small neural, autoregressive LM to greedily rescore these local predictions (subsection 7.1). While the BPD block prediction scores are produced independent of each other, neural rescoring should favor sequences that are fluent, encouraging coherence between the predictions at each head.
Parallel BPD via global n-gram rescoring   If the BPD model has $h$ heads and we consider the top-$k$ tokens from each head, then there are $k^{h}$ candidate drafts of length $h$ that can be formed. We propose to use an n-gram model to efficiently rescore all paths, using dynamic programming, and generate the $p$ most probable rescored paths as a batch of draft candidates. This draft can then be verified in parallel by the BPD model (subsection 7.2).
There are two critical distinctions between the proposed algorithms: the amount of context/expressive power available to each class of rescoring models, and fundamental limitations of decoding with each class. While neural rescoring models are potentially more expressive and can leverage unbounded context, n-gram  LMs can be used to efficiently find the globally most likely rescored drafts from the exponentially-sized set of possible draft candidates. Description of each algorithm is given in section 7.
2.3 Organization
The remainder of this paper organized as follows. In section 3, we discuss previous literature in reducing LLM latency. In section 4, we define foundational concepts and terminology. section 5 describes our experimental setup, datasets, on methods. section 6 describes our analysis of the BPD drafts. In section 7, we present the proposed BPD rescoring algorithms and empirical results, followed by a final discussion in section 8.
3 Related Work
3.1 Efficient Transformer Inference
Works on improving transformer efficiency encompass both optimization of an existing set of model weights, or a fundamental change to the model architecture. Examples of the former include techniques such as quantization (Xiao et al., 2023; Yao et al., 2022; Dettmers et al., 2022) and model pruning (Sun et al., 2023a; Ma et al., 2023). In parallel, neural architecture search has played a crucial role in identifying network structures that balance performance with efficiency (Kitaev et al., 2020; Zhu et al., 2021). Relatedly, Elbayad et al. (2019) propose early-exiting at intermediate layers for faster inference, while Schuster et al. (2022) explore confidence thresholding for balancing speed and accuracy. These methods offer insights into optimizing decoding under resource constraints.
One important line of work has focused on modifying the decoding method in LMs. The adoption of non-autoregressive (parallel) decoding strategies (Stern et al., 2018; Gu et al., 2017) marks a pivotal shift in this domain, addressing inference latency by simultaneously generating multiple tokens. Subsequent innovations have sought to refine this approach by incorporating additional context (Chi et al., 2020), iterative refinement (Kim et al., 2023), and tree-based attention mechanism (Cai et al., 2024). However, these refinements often require complex training or additional inference data.
3.2 Efficient and Effective Decoding
There are several recent works that improve the speed of LLM decoding, including pioneering works like BPD and speculative decoding. Speculative decoding leverages a smaller ‘draft’ model to anticipate the outputs of a larger target model, improving average decode latency without loss in generation quality (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023). The draft model is typically trained on the same corpus as the LLM, thus autoregressively generates similar drafts as the target model with reduced latency. Speculative decoding is most successful when a long sequence of speculated tokens are accepted by the target LM during verification, avoiding multiple serial calls to the target LM to generate the same sequence.
On the surface, contrastive decoding algorithms share some similarities with our proposed draft rescoring approach, insofar as a weaker model is used to modify the predictions of the target LM (Li et al., 2022; Kim et al., 2024). However, in this work, we refine BPD drafts solely to improve latency. Like speculative decoding, our proposals have no effect on the quality of the target LM’s generated text.
4 Preliminaries
This section introduces notation and concepts, including algorithms for standard autoregressive decoding and BPD.
Autoregressive decoding
Let $\mathcal{M}_{\theta}$ be an autoregressive LM parameterized by $\theta$. The objective is to generate an output sequence $y_{\leq T}=(y_{1},\ldots,y_{T})$ conditioned on an input sequence $\bar{x}$. $z_{t}=\mathcal{M}_{\theta}(y_{t}|\bar{x},y_{<t})$ is a vector of logits, $z_{t}\in\mathbb{R}^{|\mathcal{V}|}$, where $\mathcal{V}$ is the vocabulary over tokens. These logits define a conditional probability distribution at each time step $p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})=\frac{e^{z_{t}}}{\sum e^{z_{t}}}$, which by the chain rule yields $p_{\theta}(y_{\leq T}|\bar{x})=\prod_{t=1}^{T}p_{\theta}(y_{t}|\bar{x},y_{<t})$.
Sequences are generated autoregressively, either through ancestral sampling from some form of the conditional next token distribution (Holtzman et al., 2019), or by a beam search through the space of possible sequences to return a probable sequence. Greedy decoding, a special case of beam search, generates each token as $\hat{y}_{t+1}=\operatorname*{arg\,max}p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})$. In this work, we consider greedy decoding exclusively, as this is the setting that Stern et al. (2018) was designed to accelerate.
Algorithm 1 Blockwise Parallel Decoding (BPD)
0:  : Blockwise parallel decoder $\mathcal{M}_{\theta}^{h}$, initial prompt sequence $\bar{x}$ and target sequence length $T$.
1:  Initialize $t\leftarrow 1$
2:  while $t<T$ do
3:     /* Stage 1: Predict */
4:     $z^{1}_{t,1},\dots,z^{h}_{t,h}\leftarrow\mathcal{M}_{\theta}^{h}(y_{t+1},\dots,y_{t+h}|\bar{x},y_{\leq t})$
5:     $y_{t+1}\leftarrow\operatorname*{arg\,max}(z^{1}_{t,1})$
6:     $\hat{y}_{t+2},\dots,\hat{y}_{t+h}\leftarrow$   $\operatorname*{arg\,max}(z^{2}_{t,2}),\cdots,\operatorname*{arg\,max}(z^{h}_{t,h})$
7:     /* Stage 2: Verify */
8:     for $j\leftarrow 2,\dots,h$ in parallel do
9:        $z_{t,j}^{1},\cdots\leftarrow$  $\mathcal{M}_{\theta}^{h}(y_{t+j},\dots|\bar{x},y_{\leq{t+1}},\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1})$
10:     end for
11:     /* Stage 3: Accept */
12:     $n\leftarrow\max\{n:\hat{y}_{t+j}=\operatorname*{arg\,max}z^{1}_{t,j},2\leq j\leq n\}$
13:     $t\leftarrow t+n$
14:  end while
Blockwise parallel decoding
Let $\mathcal{M}^{h}_{\theta}$ be a BPD LM with block size $h$. This model employs $h$ distinct feedforward neural (FFN) layer with a single hidden layer, atop the target LM’s final hidden layer. The output of each FFN is followed by a softmax layer over the vocabulary to predict each of the $h$ subsequent tokens in the block. In our experiments, the parameters of the FFNs are learned jointly with the base LM during training, and the weights of all softmax layers are tied to the input embedding table. Algorithm 1 describes the BPD greedy decoding procedure:
  1. 1.
    Predict: $\mathcal{M}_{\theta}^{h}$ is used to generate a draft of $h$ token predictions $y_{t+1},\hat{y}_{t+2}\cdots,\hat{y}_{t+h}$, conditioned on the prompt, $\bar{x}$, and existing generated text, $y_{\leq t}$. $y_{t+1}$ is identical to the target LM greedy decode.
  2. 2.
    Verify: At this stage, the target LM greedily generates next-token logits $\{z_{t,2}^{1},\cdots,z_{t,h}^{1}\}$ conditioned on the existing prefix and BPD draft $\{\bar{x},y_{\leq t+1},\hat{y}_{t+2},\cdots,\hat{y}_{t+h}\}$. Verification amounts to checking which BPD draft tokens match the autoregressive greedy decode from the target LM: ($\operatorname*{arg\,max}{z^{1}_{t,i}})==\hat{y}_{t+i}$. Verification of all positions can be performed in parallel under the assumption that the target LM is a decoder-only transformer.
  3. 3.
    Accept: Finally, the length of the longest contiguous prefix of draft tokens that match the target LM greedy decode is identified: $n$. The decoded sequence is extended by $n+1$ tokens and we iterate. [^1] The decoded sequence is extended by $n+1$ tokens since during verification we generate the token from the target LM, $\operatorname*{arg\,max}{z_{t,n+1}}^{1}$, at the first position where the draft differs from the target LM greedy decode. Note that in general, not all $h$ tokens are accepted, and many of the draft tokens in each block are discarded. Since the additional time required to generate a block of tokens is fast relative to the time it takes for the forward pass of the target LM, a modest gain in accepted prefix length justifies the cost of draft generation.
5 Experimental Setup
Table 1: Per-task test performance of each finetuned model and block efficiency. Tasks include Language Modeling (LM), extractive Question Answering (QA), and both Long and Short Summarization (L-Sum & S-Sum). The performance metric for LM is perplexity, for QA is exact match, and for the remaining summarization tasks, the metric is ROUGE-L.
Task Dataset Performance Block
Efficiency
LM LAMBADA (Paperno et al., 2016) 7.88 3.12
QA SQuAD V1 (Rajpurkar et al., 2016) 57.60 2.08
S-SUM CNN/Daily (Hermann et al., 2015) 39.85 1.74
SAMSUM (Gliwa et al., 2019) 37.66 1.27
L-SUM MultiNews (Fabbri et al., 2019) 23.08 1.10
XSUM (Narayan et al., 2018) 52.15 1.13
NewsRoom (Grusky et al., 2018) 39.85 1.08
In this paper, we use $\approx 1.5$ billion (B) parameter decoder-only transformer LMs with up to nine blockwise heads. [^2] This study is based on the original BPD framework, with a modification: we use decoder-only models instead of the T5 encoder-decoder architecture. Other setups are consistent with the approach outlined in Stern et al. (2018). The 1.5B model and all other LMs were pretrained using the C4 English mixture (Raffel et al., 2019) with the causal next token prediction objective tokenized with the GPT3 subword vocabulary (Brown et al., 2020). For the 1.5B BPD models, all heads were trained jointly to predict the following $h$ tokens at each iteration. During pretraining we use batches of 2048 subword sequences, each 512 tokens in length, amounting to $\approx 200$ billion input tokens in total.
We evaluate the potential latency improvement of BPD drafts by block efficiency (Leviathan et al., 2023; Sun et al., 2023b). In this context, block efficiency represents the theoretical speedup compared to standard greedy decoding. It is defined as the average number of tokens decoded per serial call to the BPD model. The formula for block efficiency is
$B\coloneqq\frac{\text{Total number of decoded tokens}}{\text{Total number of serial calls to }\mathcal{M}_{\theta}^{h}}$
In this definition, the total number of decoded tokens is the sum of the number of accepted tokens across decoding steps, not necessarily all $h$ predicted tokens in each block. Only the tokens that pass the ‘Verify’ stage and align with the base model’s predictions are accepted and integrated into the final sequence. This ensures that generated text is identical to the target LM, while achieving speedup. The total number of serial calls to $\mathcal{M}_{\theta}^{h}$ is the number of times the model processes a block of tokens. A block efficiency of 1 means that one is achieving no speedup relative to standard decoding.
In addition to a standard language modeling dataset, LAMBADA (Paperno et al., 2016), we conduct experiments across several classes of downstream tasks. In the realm of text summarization, we evaluate models on the XSUM (Narayan et al., 2018), MultiNews (Fabbri et al., 2019), SAMSum (Gliwa et al., 2019), NewsRoom (Grusky et al., 2018) and CNN/DailyMail (Hermann et al., 2015) datasets. Each of these datasets is characterized by distinct summary lengths and styles. For extractive QA, the SQuAD V1 dataset (Rajpurkar et al., 2016) serves as our testbed. For each task aside from language modeling, we finetune the BPD model for that task. [^3] Details are given in Appendix C.
Table 1 shows that block efficiency varies dramatically across tasks and as a function of the number of BPD heads. We use all nine BPD heads for subsequent experiments as this acts as an upper bound on possible block efficiency.
Table 2: An illustration of outputs from BPD Frameworks on LAMBADA, SQuAD V1, and CNN/Daily datasets. ‘Black’ indicates standard decoded output, ‘Blue’ indicates draft tokens that were accepted during BPD, and ‘Brown’ is the prompt text.
LAMBADA
it’s nothing more than a faceless, formless brown blob to me, but I take his word for the resemblance to our genetic makeup. {Skip}
SQuAD V1
Question: Who was announced as the LEM contractor in November 1962? context: Wiesner kept up the pressure, even making the disagreement public {Skip}
Answer: Grumman
XSUM
Summarize: {Skip}
Millions of small businesses will benefit from a reduction of business rate from the Budget Osborne, Chancellor George Osborne has announced.
Table 2 sketches how BPD acts on three examples from each class of tasks.
  • LM: BPD excels at generating common multi-word expressions in a single step. For example, (no) ‘thing more than’, and (take) ‘his word for the’ are each drafted and accepted in a single step.
  • QA: BPD also attains high block efficiency in extractive QA, where it correctly drafts multi-token entities copied from the input sequence. In SQuAD V1, it accurately completes the answer ‘Grumman’ from ‘Gru’ by adding ‘mman’, highlighting its ability to process multiple tokens at once and quickly extend answers.
  • SUM: BPD’s effectiveness in SUM tasks varies by dataset. For formulaic summaries like CNN/DailyMail, it performs well, reflecting its alignment with LM and QA tasks. However, in narrative-driven datasets like SAMSum and XSUM, where concise summaries are required, the block efficiency of BPD is little better than standard decoding.
6 Exploration of Predictive Dynamics in BPD
6.1 Consecutive Repetition
Table 3: Consecutive token repetition present in BPD drafts before and after C4-trained 2-gram rescoring of the top-16 lattice. “Consec” is the percentage of consecutive identical draft tokens out of all pairs of consecutive tokens. “Max Run” is the average maximum repeated subsequence length in tokens (upper bound of nine, the number of BPD heads). Higher values correspond to more egregious repetition in drafts.
Task Dataset % Consec Max Run
Vanilla 2-gram Vanilla 2-gram
LM LAMBADA 20.0 10.7 2.2 1.8
QA SQuAD V1 75.5 67.6 6.6 6.1
S-SUM CNN/Daily 46.4 21.9 3.8 2.5
SAMSUM 29.9 20.0 3.1 2.5
L-SUM MultiNews 33.6 14.7 3.1 2.1
XSUM 24.0 9.4 2.6 1.7
NewsRoom 47.2 32.1 4.1 3.3
We observe that vanilla BPD drafts are prone to significant token repetition. This is due to the fact that each head’s prediction is independent of the others, and is a limitation shared with non-autoregressive generation in general (Gu et al., 2017). Table 3 shows the proportion of consecutive tokens in BPD drafts that are identical to each other, along with the average maximum length of repeated sequences in BPD drafts across all decode time steps. We compare these statistics before and after rescoring with a 2-gram LM: a trivial rescorer, but one that can encourage local consistency between consecutive draft tokens. Strings of repeated tokens are unnatural, and unlikely to be generated by a strong base language model. Rescoring the BPD top-$k$ lattice with even a simple language model eliminates a significant amount of repetition, reducing the percentage of consecutive repeated tokens from between 9.9% to 24.5% absolute.
6.2 Confidence Across Multiple Heads
Intuitively, predicting the probability distribution of the next $i^{\text{th}}$ token becomes harder as $i$ increases. To better understand this phenomenon, we measure the confidence of the predictions by the entropy of the probability distribution. In Figure 2, we plot the normalized histogram of entropy of each head on the LAMBADA task. From the normalized histogram, it is clear that the entropy increases as we move from first head to the last head, which agrees with our intuition that hardness of predictions increases as $i$ increases.
However, we observed that the entropy of heads does not increase monotonically for all tasks. Let ¯$\overline{\mathbb{H}}[i]$ be the average entropy of head $i$ on a particular corpus, and let $h_{\max}=\max\limits_{k}\{k:\forall i<k,\overline{\mathbb{H}}[i]\leq\overline{\mathbb{H}}[i+1]\}$ , be the index of the largest head such that the average entropy of each head increases monotonically to that point. We observed a strong correlation between $h_{\max}$ and block efficiency (Figure 3). Heads with lower entropy (indicating more confident predictions) intuitively contribute more to efficiency. Nonetheless, simply maximizing the number of low-entropy heads is not optimal, but rather incorporating progressively higher entropy heads, up to a certain point, can benefit decoding efficiency. A linear regression confirms this with an R-value of $\mathbf{0.77}$. This analysis suggests that BPD head entropy could be used as a proxy for block efficiency, and thus inference latency.
Figure 2: Entropy distributions across BPD heads on LAMBADA. The density plots illustrate the entropy distribution for each head in the model.
Figure 3: Correlation between block efficiency and $h_{\max}$, the head till which the average entropy in a task increases monotonically.
6.3 Oracle Top-$k$ Block Efficiency
Figure 4: An example of a top-5 BPD sausage lattice generated on a NewsRoom example. Edge weights correspond to (rescored) logits. Edges at each time step are ordered in descending weight and green, bolded edges correspond to candidates matching the greedy decode over the next nine tokens: ”… desktop computers with new Intel Corp processors that it …”. The initial node in this graph is state 0 and the final node is 9.
Figure 5: Illustration of the output through oracle selection. For a given top $k$ of 3, if we can choose the oracle path successfully, the block efficiency can be improved from 1 to 5.
Oracle efficiency
The concept of oracle block efficiency in BPD serves as a theoretical benchmark, illustrating the headroom available from improving the quality of the BPD draft. To compute oracle block efficiency, we consider the top-$k$ most probable tokens at each head, and form a “sausage” lattice from these. This data structure is a weighted directed graph, which succinctly represents all possible drafts (and their score under the BPD model) that could be formed from selecting one of $k$ tokens from each of the $h$ heads (Figure 4). In the automatic speech recognition and machine translation communities, it is known as a “confusion network” (Shen et al., 2006; Koehn et al., 2007).
Given the top-$k$ BPD lattice at each decoding step, we identify an oracle path that represents the path through the lattice that maximizes the length of the accepted prefix. This exercise, as shown in Figure 5, gives us insight into how much headroom exists in improving BPD drafts.
Potential headroom from oracle selection
Oracle drafting is not practical, but rather a reference point. Analyzing the gap between actual BPD performance and the oracle upper bound (Figure 6) helps us to understand the limitations of the original BPD drafts and potential areas for improvement. Additionally, exploring oracle efficiency as a function of the $k$ in the top-$k$ lattice, demonstrates how “close” the BPD draft was to producing a stronger draft.
(a) LAMBADA
(b) SQuAD V1
(c) CNN/Daily
(d) SAMSUM
(e) MultiNews
(f) XSUM
Figure 6: Oracle block efficiency over the BPD top-$k$ lattice as a function $k$. Each plot (a-f) represents a different task, demonstrating the relative improvement in block efficiency of the oracle draft with respect to the standard BPD draft as a function of the number of BPD heads used.
7 Lattice Rescoring for Improved Block Efficiency
Having explored BPD’s prediction dynamics, we propose two drafting algorithms to improve block efficiency through rescoring of the BPD lattice. This section presents techniques for rescoring the top-$k$ BPD lattice along with empirical results.
Each of these algorithms is a modification of the BPD drafted in Stage 1 in Algorithm 1. Instead of using the most likely token at each head as the prediction, we construct the top-$k$ sausage lattice of likely drafts from each head, where the set of top-$k$ tokens is denoted as $S_{i}$ for head $i$. This approach allows any token within $S_{i}$ to be chosen for position $i$, yielding a total possible combinations of:
$|S_{1}|\times|S_{2}|\times\ldots|S_{h}|=k^{h}$
In this lattice, any path from the start to final state represents a viable draft. Two algorithms are proposed to select a small number of $h$-length drafts from this lattice, which are then passed to the verification step. The first algorithm employs neural autoregressive transformers (subsection 7.1), while the second utilizes n-gram language models (subsection 7.2).
Algorithm 2 Local rescoring via neural models
0:  : Blockwise parallel decoder $\mathcal{M}_{\theta}^{h}$, top-$k$ indices selection function Top-$k(\cdot)$, rescoring model $\mathcal{M}_{\theta_{r}}$, interpolation weight $\alpha>0$.
1:  $z^{1}_{t+1},\dots,z^{h}_{t+h}\leftarrow\mathcal{M}_{\theta}^{h}(y_{t+1},\dots,y_{t+h}|\bar{x},y_{\leq t})$
2:  $S_{1},\dots,S_{h}\leftarrow\textsc{Top-}k(z^{1}_{t+1},\dots,z^{h}_{t+h})$
3:  /* Lattice Rescoring (subsection 7.1)*/
4:  for $j\leftarrow 2,\dots,h$ in parallel do
5:     $r_{t+j}\leftarrow\mathcal{M}_{\theta_{r}}(y_{t+j}|x,y_{<t+j})$
6:     $z^{j}_{t,j}[S_{j}]\leftarrow z^{j}_{t,j}[S_{j}]+\alpha\cdot r_{t+j}[S_{j}]$
7:  end for
Table 4: Block efficiency of various lattice rescoring methods over the top-$16$ BPD lattice. Parallel BPD variants consider the $16$ most probable rescored paths through the lattice. “Parallel Base BPD” indicates parallel BPD performance over the original BPD lattice. Relative percent improvement over Base is indicated in parentheses.
Task Dataset Base 4-gram neural-61M Parallel Base BPD Parallel 4-gram BPD Oracle (k=16)
[…]
7.1 Local Rescoring via Neural Models
A simple approach uses a small neural rescorer, interpolating between the logits of the rescorer LM and BPD logits with a tuned interpolation weight (line 3-6 in Algorithm 2). The locally rescored prediction is given by:
$z^{j}_{t,j}[S_{j}]\leftarrow z^{j}_{t,j}[S_{j}]+\alpha\cdot r_{t+j}[S_{j}]$
where $z^{j}_{t,j}$ represents the BPD logit vector at head $j$, and $r_{t+j}$ is the vector of logits predicted by the small neural rescoring model, which is conditioned on the sequence $y_{\leq t},\cdots,\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1}$. The parameter $\alpha$ is the weight placed on the rescorer’s prediction. We experiment with decoder-only transformers having 32, 61, and 94 million (M) weight parameters (Appendix C). We use greedy rescoring when generating the neural draft.
7.2 Parallel BPD via Global n-gram Rescoring
We also evaluate the quality of drafts generated by rescoring with an n-gram LM. Recall that BPD LMs can be used to compute a lattice representing $k^{h}$ possible sequences. We rescore all of these sequences with an n-gram model, select the top $p$ sequences and pass them to the verification stage. When $p=1$, we refer to this as n-gram rescoring and when $p>1$, we refer to this as the Parallel BPD.
While global rescoring typically yields better results compared to local rescoring, rescoring $k^{h}$ sequences with a neural LM and selecting the most likely sequence would take time $O(k^{h})$, which is computationally prohibitive in most cases. Hence, we take advantage of n-gram models, which are unique in that we can select the most likely sequence in time $\text{poly}(k,h)$ via dynamic programming. We use the OpenFST library (Allauzen et al., 2007) to represent each n-gram model as a weighted finite state automaton and apply finite state composition with the top-$k$ lattice followed by extraction of the $p$ most likely draft sequences. Training details for the n-gram models are given in Section C.3.
(a) LAMBADA
(b) SQuAD V1
(c) CNN/Daily
(d) SAMSUM
(e) MultiNews
(f) XSUM
Figure 7: Block efficiency of parallel BPD as a function of the number of top $p$ sequences verified in parallel. Each plot (a-f) displays the block efficiency achieved with parallel BPD after n-gram -rescoring compared to the the same number of paths extracted from the original top-$16$ BPD lattice. The number of top $p$ sequences is varied in ${1,2,4,8,16}$ over a top-$16$ BPD lattice.
7.3 Empirical Evaluation
Block efficiency Table 4 and Figure 7 demonstrate the impact of lattice rescoring on block efficiency across various tasks. Autoregressive neural, n-gram LM, and parallel BPD rescoring all demonstrate improvements in block efficiency, although gains are task-dependent.
  • Tasks with high initial block efficiency (LAMBADA, CNN/Daily): Both rescoring methods show little to no improvement, suggesting that vanilla BPD already produces high quality drafts.
  • Tasks with low initial block efficiency (SQuAD V1, SAMSUM, XSUM, NewsRoom): Both neural and n-gram rescoring lead to block efficiency gains, particularly with neural LMs achieving the best performance in some cases. This suggests that rescoring helps refine predictions and navigate the lattice more efficiently in these scenarios.
Table 5: Wins, ties, and losses of 61M parameter neural-rescored drafts and vanilla BPD. “Repair” corresponds to instances where the rescored draft eliminates repetition and “Regress” corresponds to instances where the rescored draft introduces repetition.
Dataset Ties Win Loss
Total % Repair % Regress Total % Repair % Regress
LAMBADA 631.5K 5804 27.95 0.05 9466 2.01 0.06
SQuAD V1 104.4K 1624 12.68 8.13 6325 2.53 12.28
CNN/Daily 965.0K 5928 23.20 0.67 17863 3.19 0.48
SAMSum 12.1K 2462 17.91 23.56 867 18.57 16.72
MultiNews 1.45M 294856 44.41 7.45 50209 22.21 5.37
XSUM 262.0K 36010 29.87 0.77 6826 4.19 10.99
NewsRoom 251.3K 79710 66.23 0.60 6492 2.85 7.39
Repairing repetitions
In subsection 6.1, we note that vanilla BPD drafts are prone to token-level repetition and that rescoring with a simple language model reduces the incidence of this. Although rescoring reduces repetition overall in drafts, is this driving improvements in block efficiency? To answer this, we compared the drafts generated by greedy rescoring with the 61M parameter neural rescorer against vanilla BPD. Time step instances were considered wins/ties/losses based on the accepted prefix length of the rescored draft vs. vanilla BPD. Table 5 displays the win frequency across tasks along with the percentage of wins/losses that can be attributed to introducing/eliminating repetition.
Note that in the tasks where rescoring improves block efficiency the most, NewsRoom and MultiNews, a high percentage of those repaired instances are driven by fixing erroneously repeated tokens. In fact, for MultiNews, 66.23% of BPD drafts are improved through repetition repair.
We also evaluated the performance of rescoring with in-domain trained rescoring LMs, but found that they tended to perform no better than C4-trained LMs (Appendix D).
8 Conclusion
In this paper, we present a comprehensive analysis of BPD, highlighting its predictive dynamics and proposing methods to enhance its inference speed. Our study offers insights into BPD’s behavior, particularly the tendency for drafts to contain consecutive repetitions and its heads to exhibit varying confidence levels in predictions. We introduce a novel measure, oracle top-$k$ block efficiency, to explore potential improvements in block efficiency. Two algorithms are proposed: one using neural models for local rescoring and another employing n-gram models for Parallel BPD. These algorithms leverage the strengths of both BPD and small rescoring models to accelerate the decoding process, pushing the boundaries of efficient text generation with BPD. We believe that this paper lays the groundwork for future explorations in optimizing language models for efficient decoding speed without compromising the output quality.
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted in the body. However, from a broader perspective, there are some potential implications related to systematic impact. These discussions are detailed further in the Appendix A.
Appendix A Broader Impact
Our research in advancing blockwise parallel decoding (BPD) for language models (LMs) paves the way for substantial improvements in language processing. This section outlines the broader implications of our work, focusing on key areas such as efficiency, scalability, and systematic impact.
  • Inference Efficiency: We’ve achieved further acceleration in inference with minimal increases in CPU overhead and memory bandwidth utilization.
  • Scalability: The introduction of FST for selecting top $p$ sequences from top $k$ predictions demonstrates notable scalability, hinting at accelerated inference with growing model sizes. This scalability is further explored through diverse experiments involving various n-gram  and neural LM architectures.
  • Systematic impact: The advancement of BPD promises identical performance to traditional autoregressive methods, but with significantly improved efficiency in decoding times. In parallel with speculative decoding research, our approach can reshape the landscape of LLM, offering faster, more accessible AI solutions across various industries.
Appendix B Limitation & Future Work
B.1 Limitation
Extending beyond greedy decoding
In this work, we focus on refining BPD drafts for greedy decoding, the decoding scheme considered in the original BPD paper (Stern et al., 2018). Extending this methodology to non-greedy sampling strategies is an area for future work. Leviathan et al. (2023) illuminates the feasibility of incorporating speculative decoding strategies to expedite non-greedy sampling processes. Their approach underscores the universality of the verification process across diverse model architectures, contingent on the availability of draft token-level predictions and probabilities — a requirement met by (rescored) BPD candidates.
Scaling up to larger models
We show that lattice rescoring can improve the block efficiency for a 1.5 billion parameter BPD LM. It remains an open question as to whether one would observe similar improvements in block efficiency with larger LMs.
Memory overhead during parallel verification
A notable challenge in our framework is the memory overhead encountered during parallel verification. The simultaneous processing of multiple token predictions, particularly for larger models and draft batches when coupled with lattice rescoring, can intensify memory demands. That said, this can be mitigated by the application of tree attention during verification (Spector & Re, 2023).
Naive drafting heads and training recipes
The implementation of drafting heads within our BPD framework, while foundational, leaves room for advanced development. This aspect is particularly crucial for refining the model’s accuracy and efficiency. Future iterations could benefit from more sophisticated training methods, enhancing these heads’ ability to navigate complex linguistic contexts and improving overall predictive performance.
B.2 Future Work
The evolution of the BPD framework will be pivotal in addressing its current constraints and broadening its utility. Key areas for future research include:
  • The intrinsic compatibility of our lattice rescoring method with alternative sampling strategies presents a fertile ground for future inquiries. This delineation not only enriches the discourse within efficient language model inference but also sets a trajectory for subsequent empirical endeavors.
  • Scaling the BPD model for compatibility with larger-scale LLMs, ensuring it remains effective and efficient as language models evolve.
  • Advancing training methodologies for drafting heads, to bolster their predictive accuracy and contribution to the BPD process, thereby optimizing the framework for more complex language modeling tasks.
  • Using the sequential entropy ordering of heads (Figure 3) as a possible halting condition during BPD head training, or to inform how a rescoring LM should be interpolated with the BPD lattice weights.
Appendix C Experiment Details
C.1 Training Objective for BPD LMs
We minimized the following loss function to train BPD LMs:
$\mathcal{L}_{BPD}=\sum_{h=1}^{H}\lambda_{h}\mathcal{L}_{h}$
where $H$ is the number of heads, $\lambda_{h}$ is a non-negative scalar that weights the loss from head $h$, and $\mathcal{L}_{h}$ denotes the loss for each individual head:
$\mathcal{L}_{h}=-\sum_{x_{1\ldots i},y_{i+h}}\log p(y_{i+h}|x_{1\ldots i})$
where $x_{1\ldots i}$ is the token sequence up to position $i$, $y_{i+h}$ is the ground truth token at position $i+h$, and $p(y_{i+h}|x_{1\ldots i})$ is the probability of observing token $y_{i+h}$ given the sequence $x_{1\ldots i}$ under the BPD LM. We trained all models in this work with $\lambda_{h}=1$. We leave tuning these hyperparameters, improving the block efficiency and quality of the BPD LM, as future work.
C.2 Neural Model Details
Table 6: Architecture hyperparameters for each of the transformer-based neural language models.
Type Model # Layers Embedding Dim Hidden Dim
Blockwise Parallel Decoder 1.5B 18 1,536 12,288
Autoregressive Decoder 32M 2 384 1,536
61M 12 384 1,536
94M 6 768 3,072
Each neural rescoring LM is a decoder-only transformer with learned absolute positional embeddings and twelve self-attention heads at each layer. The key architecture hyperparameters are given in Table 6. Aside from scale, the only difference between the BPD model and neural rescoring models is the addition of the feedforward neural networks and eight additional block prediction heads. Note that the number of parameters for each of these models also includes the embedding table.
Each model was pretrained on the English C4 corpus for 200K iterations with a batch size of $2^{20}\approx 1M$ tokens per batch. Dropout was not applied. For the BPD model, all heads were trained jointly.
For downstream tasks, models were finetuned for a maximum 100K iterations with a batch size of two examples with maximum sequence length of 2048. Maximum learning rate was fixed to $10^{-4}$ for all runs, with a cosine learning rate schedule. Checkpoints were selected based on heldout set model performance. Interpolation weight for all rescoring models was tuned for block efficiency on 100 randomly selected examples from the evaluation set for each task, and performance was reported on the remainder of the evaluation set.
C.3 n-gram Details
All n-gram  LMs in this work are Katz backoff n-gram LMs (Katz, 1987) fit on the train split of the GPT3 subword-tokenized English C4 corpus with n-gram order $\in\{2,4\}$. We apply entropy pruning (Stolcke, 2000) to reduce model size to a maximum of 100 million n-grams per model, and ensure that each trigram is observed at least twice and each 4-gram is observed at least four times. Preprocessing of the text is identical to that used to train neural LMs.
C.4 Datasets
  • LAMBADA (LAnguage Modeling Broadened to Account for Discourse Aspects): A collection of narrative passages designed to test the understanding of long-range dependencies in language models, where the task involves predicting the last word of a passage based on the full context (Paperno et al., 2016).
  • SQuAD V1 (Stanford Question Answering Dataset): A reading comprehension dataset that features questions based on Wikipedia articles, with answers located within the text (Rajpurkar et al., 2016).
  • CNN/DailyMail: This dataset includes news articles paired with human-written summaries, mainly used to evaluate the summarization capabilities of language models, particularly in abstractive summarization (Hermann et al., 2015).
  • SAMSum (Semi-Automatic Machine Summarization): Focuses on abstractive summarization using news articles and machine-generated summaries, testing models’ abilities to refine and improve existing summaries (Gliwa et al., 2019).
  • MultiNews: Comprises news articles from diverse sources for abstractive summarization tasks, evaluating models on handling different writing styles and topics (Fabbri et al., 2019).
  • XSUM: Contains scientific documents and summaries, challenging language models to process complex scientific information and language (Narayan et al., 2018).
  • NewsRoom: A dataset of news articles aimed at assessing the factual accuracy and information extraction capabilities of models in generating summaries (Grusky et al., 2018).
All datasets were tokenized using the 50,257 GPT3 subword vocabulary (Brown et al., 2020).
Templates
We used the following prompts during model finetuning and inference.
  • SQuAD: ”question: [question] context: [context]”
  • CNN/DailyMail: ”summarize: [text]”
[…]
  • MultiNews: ”Write a summary based on this article: [text]”
  • XSUM: ”Summarize: [text]”
  • NewsRoom: ”Please write a short summary for the following article: [title] [text]”
Appendix D Rescoring with In-domain Language Models
Table 7: Block efficiency from rescoring with in-domain trained rescoring models for 2-gram and 61M parameter neural rescorer.
Dataset 2-gram neural-61M
C4 In-domain C4 In-domain
SQuAD V1 2.09 2.04 2.10 2.06
CNN/Daily 1.73 1.73 1.73 1.72
SAMSUM 1.31 1.22 1.39 1.24
MultiNews 1.13 1.14 1.25 1.16
XSUM 1.17 1.18 1.23 1.14
NewsRoom 1.20 1.22 1.29 1.11
We found that in-domain rescorers performed no better than rescorers only trained on C4. We suspect this is due to a lack of sufficient finetuning data and that the main benefit of rescoring comes from discouraging unnatural artifacts such as repetition from the original BPD draft. Table 7 shows block efficiency after rescoring using in-domain models for all tasks besides language modeling.
Neural rescorers were finetuned from C4-pretrained checkpoints. n-gram  models were trained from scratch, and unseen vocabulary was added as unigram arcs with trivial weight (negative log probability of 1000.0). This was done to ensure that all paths through the lattice were assigned non-zero probability by the n-gram  model. In previous experiments, we also tried interpolating the in-domain n-gram  model with a unigram model trained on C4, and observed similar performance as simply adding unseen unigrams.
Appendix E Interpolation Weights Tuned per Task
We tuned the interpolation weight, $\alpha$ for the 94M parameter neural and 4-gram LM rescorers, and then used this weight to rescore with all other models of that same class. 100 examples from each task’s heldout set were set aside for tuning, to maximize block efficiency. The remainder of examples were used for evaluation. We swept over $\alpha\in\{0.1,0.5,0.75,0.9,1.0,1.1,1.5,2.0,5.0,10.0\}$.
Note that for tasks where lattice rescoring was unhelpful, the interpolation weight, $\alpha$ is tuned to place much higher weight on the BPD logits (Table 8). This is a signal that the rescorer does not provide additional information over the original BPD heads.
Table 8: Tuned interpolation weight per task for neural and n-gram rescoring.
Dataset Neural n-gram
LAMBADA 0.1 0.1
SQuAD V1 1.0 0.75
SAMSum 5.0 1.5
CNN/Daily 0.1 0.1
MultiNews 5.0 2.0
XSUM 1.5 1.1
NewsRoom 5.0 2.0
Appendix F Local Rescoring Impact on Block Efficiency
[…]
Task Dataset Base 2-gram 3-gram 4-gram neural-32M neural-61M neural-94M Oracle (k=2) Oracle (k=16)
[…]
Table 9 reveals the impact of different rescoring methods on the block efficiency of the BPD lattice, offering valuable insights into their effectiveness across diverse tasks and models, supporting the investigations in section 7.
  • Limited improvement for high baselines: For tasks with already high initial block efficiency (LAMBADA, CNN/DailyMail), rescoring offers minimal or even negative changes in block efficiency compared to the baseline BPD system. This suggests that for well-calibrated models with hierarchical confidence ordering, the standard BPD already achieves significant speed improvements, leaving limited room for further gains through rescoring.
  • Efficacy for poor baselines: In tasks with less calibrated predictions (SQuAD V1, XSUM, NewsRoom), rescoring using both N-gram and neural language models demonstrably improves block efficiency. Notably, neural rescoring with larger models (61M and 94M parameters) achieves the highest efficiency gains in these tasks, reaching up to 19.44% improvement in NewsRoom. These results highlight the potential of rescoring to refine predictions and enhance efficiency for models exhibiting calibration issues.
  • Task-specific effectiveness: The level of improvement from rescoring varies across different summarization tasks (MultiNews, XSUM, NewsRoom). While all show positive responses, NewsRoom exhibits the most significant gains, suggesting that the effectiveness of rescoring can be task-dependent.
  • Comparison with oracle efficiency: The ‘Oracle’ columns present the upper bound achievable if only the most likely token at each step is chosen with perfect hindsight (k=2 and k=16). While significant gaps remain between current results and the oracle, the observed improvements from rescoring demonstrate progress towards closing this efficiency gap.
Overall, these findings suggest that local rescoring methods can be a valuable tool for enhancing BPD efficiency, particularly for models with less calibrated predictions. Further exploration of advanced rescoring strategies, especially in conjunction with larger neural language models, holds promise for achieving even closer-to-oracle efficiency levels.
Appendix G Local Rescoring via Neural Models vs Parallel BPD via Global n-gram Rescoring
This section supports the detailed explanations on two algorithms in section 7: local rescoring via neural models and parallel BPD via global n-gram rescoring.
  • Local rescoring via neural models typically refers to a process where predictions are adjusted at each individual step of the sequence generation. A local rescoring algorithm takes the initial predictions from the model and refine them using additional information or a secondary model.
  • Parallel BPD via global n-gram rescoring, on the other hand, implies a more comprehensive approach where the entire sequence or multiple sequences are considered together. The ‘global’ aspect means that the rescoring process takes into account the whole sequence’s context to find the optimal path, rather than making decisions based on individual tokens. Parallel batched processing indicates that multiple sequences are being evaluated and rescored in one go, which is often more efficient. n-gram rescoring uses the statistical properties of n-grams to predict the probability of a sequence, allowing for the consideration of the sequence as a whole rather than as isolated predictions.
Appendix H Ablation on the Number of Heads in the BPD Model
Table 10 summarizes the block efficiency for different head configurations across various language tasks with the same settings discussed in Table 1.
  • General trend: Both performance and block efficiency tend to increase with the number of heads, up to a point. This suggests that using more heads allows the model to capture richer contextual information and make more accurate predictions.
  • Efficiency trade-off: While increasing heads generally improves block efficiency, it also increases the memory for verification stages. Therefore, the optimal number of heads depends on the balance between desired block efficiency and available resources.
Table 10: Test performance per task. Test performance of each finetuned model and block efficiency are shown as a function of heads ($h\in{3,6,9}$). Tasks inclue Language Modeling (LM), extractive Question Answering (QA), and both Long and Short Summarization (L-Sum & S-Sum). The metric for LM is perplexity, for QA is exact match, and for all the remaining (summarization) tasks, the metric is ROUGE-L.
Task Dataset Performance # of Heads ($h$)
3 6 9
LM LAMBADA 7.88 1.79 2.84 3.12
QA SQuAD V1 57.60 1.53 2.03 2.08
S-SUM CNN/Daily 39.85 1.60 1.71 1.74
SAMSUM 37.66 1.18 1.25 1.27
L-SUM MultiNews 23.08 1.08 1.08 1.10
XSUM 52.15 1.11 1.12 1.13
NewsRoom 39.85 1.07 1.08 1.08
Appendix I Efficiency of BPD Rescoring
To enhance our understanding of BPD rescoring within the realm of contemporary deep learning hardware environments, we present an in-depth examination focused on TPU/GPU utilization and the overhead incurred by n-gram  rescoring. This analysis is divided into two parts: (1) an analysis of BPD rescoring through the lens of TPU/GPU utilization, and (2) empirical benchmarks of n-gram  lattice rescoring. The major takeaways are as follows.
Memory bandwidth (HBM $\Leftrightarrow$ SRAM)
A critical factor in the performance of deep learning applications is the efficient management of memory bandwidth between High Bandwidth Memory (HBM) and Static Random Access Memory (SRAM) (Dao et al., 2022). Increasing the block efficiency via BPD lattice rescoring reduces the average per token parameter and key-value cache I/O that needs to be communicated from HBM to SRAM.
Overhead in n-gram  rescoring
n-gram  rescoring is actually quite efficient. For the size of lattices we consider in this work, moving the lattice from HBM to DRAM, performing n-best n-gram  rescoring, and moving the n-best paths back to HBM requires no more than 2 ms per lattice.
I.1 TPU/GPU Utilization
We compare our approach against traditional Autoregressive LMs across several metrics (Table 11):
Table 11: Comparative analysis of per decoded token efficiency metrics across BPD rescoring methods and the traditional autoregressive LM (batch size=1). This table showcases the average block efficiency, parameter I/O, key-value (KV) cache I/O at varying sequence lengths, and FLOPS—evaluated on a per-token basis under the condition of batch size 1.
Component Autoregressive Base BPD 4-gram BPD Neural-61M BPD Parallel Base BPD Parallel 4-gram BPD
Avg. Block Efficiency 1.000 1.646 1.657 1.724 1.717 1.797
Parameter I/O (GB) 3.000 1.823 1.811 1.811 1.747 1.669
KV Cache I/O (GB) - Seq_len 128 0.113 0.074 0.073 0.076 0.140 0.134
KV Cache I/O (GB) - Seq_len 512 0.453 0.280 0.278 0.290 0.338 0.323
KV Cache I/O (GB) - Seq_len 1024 0.906 0.555 0.552 0.574 0.602 0.575
KV Cache I/O (GB) - Seq_len 2048 1.812 1.106 1.098 1.144 1.129 1.079
FLOPS (T) 0.931 0.57 0.567 0.635 0.621 0.593
Memory bandwidth and compute efficiency
The BPD rescoring variants achieve significant reductions in Parameter I/O and KV Cache I/O compared to autoregressive decoding, suggesting BPD methods’ ability to reducing inference times by mitigating the primary latency bottleneck—memory bandwidth. Advances in TPU/GPU architecture ensure that an increase in FLOPS per token has a minimal effect on latency, confirming our strategy’s capacity to navigate the complexities of memory bandwidth efficiently.
Comparative latency impact
A consistent decrease in memory bandwidth utilization across BPD LMs, including those leveraging LM rescoring and parallel processing strategies, illustrates our approach’s contribution to accelerating inference speed. This underscores the practicality and applicability of our enhancements in promoting more efficient language model inference within state-of-the-art computational frameworks.
I.2 Overhead of n-gram Rescoring
While the majority of computational efforts in BPD rescoring are dedicated to TPU/GPU utilization, the implementation of n-gram rescoring introduces additional overheads. These are primarily attributed to CPU computations and the data transfer between the CPU and HBM. This section provides a comprehensive examination of these overheads, drawing on benchmarks from rescoring experiments with a 4-gram C4 LM.
Benchmarks for 4-gram C4 LM rescoring
We conducted benchmarks on rescoring lattices with a 4-gram C4 LM of $\approx$100M n-grams. The average latency observed across 10 runs for different numbers of the shortest paths is summarized in the Table 12:
Table 12: Average latency for N-best rescoring an 8-time step lattice with 16 arcs per time step. N, the number of shortest paths, is varied from 1 to 16.
# Shortest Paths N-best Rescoring Latency (ms)
1 1.630
2 1.751
4 1.878
8 1.871
16 1.983
Notably, rescoring with a large 4-gram LM averages less than 2 milliseconds for extracting up to 16 globally-best paths, despite the lattice containing approximately 4.29 billion possible paths. In our initial experiments, increasing the size of the n-gram  LM had little effect on n-best rescoring latency, indicating that enhancements leading to higher block efficiency will incur little additional latency, provided the LM fits within DRAM.
Latency is predominantly influenced by lattice size, particularly the number of top-k tokens per time step and the number of time steps, as depicted in Table 13:
Table 13: 1-best rescoring latency by the 4-gram C4 LM for varying lattice sizes.
Number of time steps Top-k per time step 1-best rescoring latency (ms)
4 2 1.038
4 4 1.050
4 8 1.130
4 16 1.237
8 2 1.061
8 4 1.144
8 8 1.234
8 16 1.630
16 2 1.102
16 4 1.206
16 8 1.558
16 16 2.174
The benchmarks highlight the fact that the additional overhead introduced by n-gram  rescoring, though present, should not significantly impact overall latency.