### v2

# Taehyeon Kim1, Ananda Theertha Suresh2 Kishore Papineni2 Michael Riley2 Sanjiv Kumar2 Adrian Benton2 1KAIST AI  2Google Research

Abstract
Despite the remarkable strides made by autoregressive language models, their potential is often hampered by the slow inference speeds inherent in sequential token generation. Blockwise parallel decoding (BPD) was proposed by Stern et al. [38] as a method to improve inference speed of language models by simultaneously predicting multiple future tokens, termed block drafts, which are subsequently verified and conditionally accepted by the autoregressive model. This paper contributes to the understanding and improvement of block drafts in two ways. First, we analyze the token distributions produced by multiple prediction heads. Secondly, we leverage this analysis to develop algorithms to improve BPD inference speed by refining the block drafts using n-gram and neural language models. Experiments demonstrate that refined block drafts yield a +5-21% increase in block efficiency (i.e., the number of accepted tokens from the block draft) across diverse datasets.
1 Introduction
The landscape of natural language processing has been profoundly reshaped by recent advances in autoregressive language models [3, 43, 30, 33, 42]. These models have shown remarkable proficiency across a range of text generation tasks, including applications like question answering [34] and summarization [15]. However, a significant obstacle to their wider application is high inference latency, particularly for extremely deep models with hundreds of billions of parameters [16, 31, 7]. This latency, intrinsic to decoding with autoregressive language models (LMs), imposes considerable computational burdens and limits real-time deployment.
In response to these latency challenges, the field has seen a shift towards decoding methods aimed at reducing the inference latency in large language models (LLM). One promising development is the concept of blockwise parallel decoding (BPD) [38, 27, 4]. Unlike autoregressive decoding, which generates one token at a time, blockwise parallel LMs are outfitted with a set of prediction heads that propose and verify a draft, a block of subsequent tokens, in parallel. While BPD offers one solution to accelerated text generation, it also poses a challenge in ensuring that the proposed drafts are fluent and natural.
(a) Example of block drafts
(b) Output of our proposed algorithms
Figure 1: (a) Illustration of two tokens that are decoded by autoregressive decoding vs. two tokens drafted by BPD. (b) Outputs from our proposed algorithms, where the top-$k$ token-level predictions are refined using local neural and global n-gram rescoring, which selects the $p$ most probable sequences by dynamic programming, for batched verification.
BPD inference speed depends on both the time it takes to produce a block draft and the draft’s agreement with the base LM’s output (1(a)). Unlike standard autoregressive LMs that generate tokens sequentially—ensuring consistency with all preceding tokens (e.g., ‘Messi’ following ‘Lionel’)—BPD employs a parallel strategy. Here, blockwise parallel LMs simultaneously predict multiple token drafts (e.g., ‘Lionel’ and ‘Ronaldo’), each independently. The primary challenge in BPD is ensuring that these concurrently generated tokens maintain consistency. Effective block drafters should prioritize coherent sequences such as ‘Lionel Messi’ over less likely combinations like ‘Lionel Ronaldo’, which a robust LM would not decode. The focus of this research is improving the quality of block drafts without altering the underlying model parameters.
2 Our contributions
In this paper, we first investigate properties made by the prediction heads of blockwise parallel LMs across several tasks; given these observations, we propose rescoring algorithms to produce higher quality block drafts.
2.1 Observations on block drafts
Consecutive repetitions  All heads within a block make predictions independently in a blockwise parallel LM. Unsurprisingly, we observe that this leads to block drafts with significant token repetition across heads. Consecutive repetition is pervasive across tasks, ranging from 20% to 75% of all neighboring draft tokens, depending on the task (subsection 6.1).
Confidence of different heads  We analyze the distribution of probabilities within each softmax head. Our empirical analysis reveals a key property of BPD: the block drafter tends to be more confident with initial tokens, and becomes progressively less confident for subsequent tokens. We find that the confidence of block heads correlates strongly with the quality of the block drafter (subsection 6.2).
Oracle top-$k$ block efficiency  In the standard BPD algorithm (Algorithm 1), the most likely token at each head is generated as the draft. As highlighted before, this approach is prone to two issues: (1) there might be consecutive repetitions and (2) the model might not be confident about the prediction at some of the heads. We use block efficiency, the average number of draft tokens accepted during decoding, to measure the quality of a given drafter [24, 41]. We ask if the block efficiency can be improved by considering the top-$k$ most likely tokens at each head. To measure the potential benefit of considering top-$k$ tokens, we measure the block efficiency of the oracle path through this top-$k$ lattice, oracle top-$k$ block efficiency, and show that there is significant headroom for improvement across tasks (subsection 6.3).
2.2 New algorithms
Based on these observations, we propose two algorithms to leverage the top-$k$ predictions at each head and improve BPD latency (1(b)). Neither of these algorithms require changes to the underlying blockwise parallel LMs.
Local rescoring via neural LMs  Given the top-$k$ predictions at each head, we refine the block draft by using a small neural, autoregressive LM to greedily rescore these local predictions (subsection 7.1). While the block prediction scores are produced independent of each other, neural rescoring should favor sequences that are fluent, encouraging coherence between the predictions at each head.
Global rescoring via n-gram LMs with multi-drafts   If the blockwise parallel LM has $h$ heads and we consider the top-$k$ tokens from each head, then there are $k^{h}$ candidate drafts of length $h$ that can be formed. We propose to use an n-gram model to efficiently rescore all paths, via dynamic programming, and generate the $p$ most probable rescored paths as a batch of draft candidates. These $p$ drafts can then be verified in parallel by the blockwise parallel LM (subsection 7.2).
Figure 2: Relative performance of our methods to standard BPD with a 1.5B parameter blockwise parallel LM on NewsRoom dataset [13]. Details are described in Appendix H.
There are two critical distinctions between the proposed algorithms: the amount of context/expressive power available to each class of rescoring model, and fundamental limitations of decoding with each class. While neural rescoring models are potentially more expressive and can leverage unbounded context, n-gram  LMs can be used to efficiently find the globally most likely rescored drafts from the exponentially-sized set of possible draft candidates. Figure 2 shows that our proposed methods enhance block efficiency, with one approach increasing it by up to +21.30%. This same method also optimizes resource usage, reducing key-value (KV) cache I/O by -2.54% and additionally using FLOPs per token by +4.04%. Description of each algorithm is given in subsection 7.1.
2.3 Organization
The remainder of this paper organized as follows. In section 3, we discuss previous literature in reducing LLM latency. In section 4, we define foundational concepts and terminology. section 5 describes our experimental setup, datasets, on methods. section 6 describes our analysis of the block drafts. In subsection 7.1, we present the proposed BPD rescoring algorithms and empirical results, followed by a final discussion in section 8.
3 Related Work
3.1 Efficient transformer inference
Works on improving transformer efficiency encompass both optimization of an existing set of model weights, or a fundamental change to the model architecture. Examples of the former include techniques such as quantization [44, 45, 9] and model pruning [40, 26]. In parallel, neural architecture search has played a crucial role in identifying network structures that balance performance with efficiency [22, 46]. Relatedly, Elbayad et al. [10] propose early-exiting at intermediate layers for faster inference, while Schuster et al. [35] explore confidence thresholding for balancing speed and accuracy. These methods offer insights into optimizing decoding under resource constraints.
One important line of work has focused on modifying the decoding method in LMs. The adoption of non-autoregressive (parallel) decoding strategies [38, 14] marks a pivotal shift in this domain, addressing inference latency by simultaneously generating multiple tokens. Subsequent innovations have sought to refine this approach by incorporating additional context [6], iterative refinement [20], and tree-based attention mechanism [4]. However, these refinements often require complex training or additional inference data.
3.2 Efficient and effective decoding
There are several recent works that improve the speed of LLM decoding, including pioneering works like BPD and speculative decoding. Speculative decoding leverages a smaller ‘draft’ model to anticipate the outputs of a larger target model, improving average decode latency without loss in generation quality [24, 5, 20]. The draft model is typically trained on the same corpus as the LLM, thus autoregressively generates similar drafts as the target model with reduced latency. Speculative decoding is most successful when a long sequence of speculated tokens are accepted by the target LM during verification, avoiding multiple serial calls to the target LM to generate the same sequence.
On the surface, contrastive decoding algorithms share some similarities with our proposed draft rescoring approach, insofar as a weaker model is used to modify the predictions of the target LM [25, 21]. However, in this work, we refine block drafts solely to improve latency. Like speculative decoding, our proposals have no effect on the quality of the target LM’s generated text.
4 Preliminaries
This section introduces notation and concepts, including algorithms for standard autoregressive decoding and BPD.
Autoregressive decoding  Let $\mathcal{M}_{\theta}$ be an autoregressive LM parameterized by $\theta$. The objective is to generate an output sequence $y_{\leq T}=(y_{1},\ldots,y_{T})$ conditioned on an input sequence $\bar{x}$. $z_{t}=\mathcal{M}_{\theta}(y_{t}|\bar{x},y_{<t})$ is a vector of logits, $z_{t}\in\mathbb{R}^{|\mathcal{V}|}$, where $\mathcal{V}$ is the vocabulary over tokens. These logits define a conditional probability distribution at each time step $p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})=\frac{e^{z_{t}}}{\sum e^{z_{t}}}$, which by the chain rule yields $p_{\theta}(y_{\leq T}|\bar{x})=\prod_{t=1}^{T}p_{\theta}(y_{t}|\bar{x},y_{<t})$.
Sequences are generated autoregressively, either through ancestral sampling from some form of the conditional next token distribution [17], or by a beam search through the space of possible sequences to return a probable sequence. Greedy decoding, a special case of beam search, generates each token as $\hat{y}_{t+1}=\operatorname*{arg\,max}p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})$. In this work, we consider greedy decoding exclusively, as this is the setting that Stern et al. [38] was designed to accelerate.
10:  : Blockwise parallel LM $\mathcal{M}_{\theta}^{h}$, initial prompt sequence $\bar{x}$ and target sequence length $T$.
21:  Initialize $t\leftarrow 1$
2:  while $t<T$ do
3:     /* Stage 1: Predict */
4:     $z^{1}_{t,1},\dots,z^{h}_{t,h}\leftarrow\mathcal{M}_{\theta}^{h}(y_{t+1},\dots,y_{t+h}|\bar{x},y_{\leq t})$
5:     $y_{t+1},\hat{y}_{t+2},\dots,\hat{y}_{t+h}\leftarrow\operatorname*{arg\,max}(z^{1}_{t,1}),\operatorname*{arg\,max}(z^{2}_{t,2}),\cdots,\operatorname*{arg\,max}(z^{h}_{t,h})$
36:     /* Stage 2: Verify */
7:     for $j\leftarrow 2,\dots,h$ in parallel do
8:        $z_{t,j}^{1},\cdots\leftarrow$$\mathcal{M}_{\theta}^{h}(y_{t+j},\dots|\bar{x},y_{\leq{t+1}},\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1})$
9:     end for
410:     /* Stage 3: Accept */
511:     $n\leftarrow\max\{n:\hat{y}_{t+j}=\operatorname*{arg\,max}z^{1}_{t,j},2\leq j\leq n\}$
612:     $t\leftarrow t+n$
13:  end while
Algorithm 1 Blockwise parallel decoding (BPD)
Blockwise parallel decoding  Let $\mathcal{M}^{h}_{\theta}$ be a blockwise parallel LM with block size $h$. This model employs $h$ distinct feedforward neural (FFN) layer with a single hidden layer, atop the target LM’s final hidden layer. The output of each FFN is followed by a softmax layer over the vocabulary to predict each of the $h$ subsequent tokens in the block. In our experiments, the parameters of the FFNs are learned jointly with the base LM during training, and the weights of all softmax layers are tied to the input embedding table. Algorithm 1 describes the BPD greedy decoding procedure:
  1. 1.
    Predict: $\mathcal{M}_{\theta}^{h}$ is used to generate a draft of $h$ token predictions $y_{t+1},\hat{y}_{t+2}\cdots,\hat{y}_{t+h}$, conditioned on the prompt, $\bar{x}$, and existing generated text, $y_{\leq t}$. $y_{t+1}$ is identical to the target LM greedy decode.
  2. 2.
    Verify: At this stage, the target LM greedily generates next-token logits $\{z_{t,2}^{1},\cdots,z_{t,h}^{1}\}$ conditioned on the existing prefix and block draft $\{\bar{x},y_{\leq t+1},\hat{y}_{t+2},\cdots,\hat{y}_{t+h}\}$. Verification amounts to checking which block draft tokens match the autoregressive greedy decode from the target LM: ($\operatorname*{arg\,max}{z^{1}_{t,i}})==\hat{y}_{t+i}$. Verification of all positions can be performed in parallel under the assumption that the target LM is a decoder-only transformer.
  3. 3.
    Accept: Finally, the length of the longest contiguous prefix of draft tokens that match the target LM greedy decode is identified: $n$. The decoded sequence is extended by $n+1$ tokens and we iterate. [^1] The decoded sequence is extended by $n+1$ tokens since during verification we generate the token from the target LM, $\operatorname*{arg\,max}{z^{1}_{t,n+1}}$, at the first position where the draft differs from the target LM greedy decode. Note that in general, not all $h$ tokens are accepted, and many of the draft tokens in each block are discarded. Since the additional time required to generate a block of tokens is fast relative to the time it takes for the forward pass of the target LM, a modest gain in accepted prefix length justifies the cost of draft generation.
5 Experimental setup
Table 1: Per-task test performance of each finetuned model and block efficiency over language modeling (LM), extractive question answering (QA), and both long and short summarization (L-Sum & S-Sum).
Task Dataset Performance Block
Efficiency
LM LAMBADA [29] 7.88 3.12
QA SQuAD V1 [34] 57.60 2.08
S-SUM CNN/Daily [15] 39.85 1.74
SAMSUM [12] 37.66 1.27
L-SUM MultiNews [11] 23.08 1.10
XSUM [28] 52.15 1.13
NewsRoom [13] 39.85 1.08
Table 2: Outputs from BPD frameworks. Black indicates standard decoded output, blue indicates accepted draft tokens, and brown is the prompt.
LAMBADA
it’s nothing more than a faceless, formless brown blob to me, but I take his word for the resemblance to our genetic makeup. {Skip}
SQuAD V1
Question: Who was announced as the LEM contractor in November 1962? context: Wiesner kept up the pressure, even making the disagreement public {Skip}
Answer: Grumman
XSUM
Summarize: {Skip}
Millions of small businesses will benefit from a reduction of business rate from the Budget Osborne, Chancellor George Osborne has announced.
In this paper, we use $\approx 1.5$ billion (B) parameter decoder-only transformer LMs with up to 9 blockwise heads. [^2] This study is based on the original BPD framework, with a modification: we use decoder-only models instead of the T5 encoder-decoder architecture. Other setups are consistent with the approach in Stern et al. [38]. The 1.5B model and all other LMs were pretrained on (English) C4 [32] with the causal next token prediction objective tokenized with the GPT3 subword vocabulary [3]. For the 1.5B blockwise parallel LMs, all heads were trained jointly to predict the following $h$ tokens at each iteration. During pretraining, we use batches of 2048 subword sequences, each 512 tokens in length, amounting to $\approx 200$B input tokens in total on TPUv3/TPUv4 [18] with Jax [2].
We evaluate the potential latency improvement of block drafts by block efficiency [24, 41]. In this context, block efficiency represents the theoretical speedup compared to standard greedy decoding. It is defined as the average number of tokens decoded per serial call to the blockwise parallel LMs. The formula for block efficiency is given by $B\coloneqq\frac{\text{Total number of decoded tokens}}{\text{Total number of serial calls to }\mathcal{M}_{\theta}^{h}}$.
In this definition, the total number of decoded tokens is the sum of the number of accepted tokens across decoding steps, not necessarily all $h$ predicted tokens in each block. Only the tokens that pass the ‘Verify’ stage and align with the base model’s predictions are accepted and integrated into the final sequence. This ensures that generated text is identical to the target LM, while achieving speedup. The total number of serial calls to $\mathcal{M}_{\theta}^{h}$ is the number of times the model processes a block of tokens. A block efficiency of 1 means that one is achieving no speedup relative to standard decoding.
In addition to a standard language modeling dataset, LAMBADA [29], we conduct experiments across several classes of downstream tasks. In the realm of text summarization, we evaluate models on the XSUM [28], MultiNews [11], SAMSum [12], NewsRoom [13] and CNN/DailyMail [15] datasets. Each of these datasets is characterized by distinct summary lengths and styles. For extractive QA, the SQuAD V1 dataset [34] serves as our testbed. For each task aside from language modeling, we finetune the blockwise parallel LM for that task. [^3] Details are given in Appendix C. Table 1444The performance metric for LM is perplexity, for QA is exact match, and for the remaining summarization tasks, the metric is ROUGE-L. shows that block efficiency varies dramatically across tasks and as a function of the number of block draft heads. We use all 9 block draft heads for subsequent experiments as this acts as an upper bound on possible block efficiency.
Table 2 sketches how BPD acts on three examples from each class of tasks.
  • LM: BPD excels at generating common multi-word expressions in a single step. For example, (no) ‘thing more than’, and (take) ‘his word for the’ are each drafted and accepted in a single step.
  • QA: BPD also attains high block efficiency in extractive QA, where it correctly drafts multi-token entities copied from the input sequence. In SQuAD V1, it accurately completes the answer ‘Grumman’ from ‘Gru’ by adding ‘mman’, highlighting its ability to process multiple tokens at once and quickly extend answers.
  • SUM: BPD’s effectiveness in SUM tasks varies by dataset. For formulaic summaries like CNN/DailyMail, it performs well, reflecting its alignment with LM and QA tasks. However, in narrative-driven datasets like SAMSum and XSUM, where concise summaries are required, the block efficiency of BPD is little better than standard decoding.
6 Exploration of predictive dynamics in BPD
6.1 Consecutive repetition
Table 3: Consecutive token repetition in block drafts before and after C4-trained 2-gram rescoring of the top-16 lattice. “Consec" is the percentage of consecutive identical draft tokens out of all pairs of consecutive tokens. “Max run" is the average maximum repeated subsequence length in tokens (upper bound of 9, the number of block draft heads). Higher values correspond to more egregious repetition in drafts.
Task Dataset % Consec Max run
Vanilla 2-gram Vanilla 2-gram
LM LAMBADA 20.0 10.7 2.2 1.8
QA SQuAD V1 75.5 67.6 6.6 6.1
S-SUM CNN/Daily 46.4 21.9 3.8 2.5
SAMSUM 29.9 20.0 3.1 2.5
L-SUM MultiNews 33.6 14.7 3.1 2.1
XSUM 24.0 9.4 2.6 1.7
NewsRoom 47.2 32.1 4.1 3.3
We observe that vanilla block drafts are prone to significant token repetition. This is due to the fact that each head’s prediction is independent of the others, and is a limitation shared with non-autoregressive generation in general [14]. Table 3 shows the proportion of consecutive tokens in block drafts that are identical to each other, along with the average maximum length of repeated sequences in block drafts across all decode time steps. We compare these statistics before and after rescoring with a 2-gram LM: a trivial rescorer, but one that can encourage local consistency between consecutive draft tokens. Strings of repeated tokens are unnatural, and unlikely to be generated by a strong base language model. Rescoring the top-$k$ lattice with even a simple language model eliminates a significant amount of repetition, reducing the percentage of consecutive repeated tokens from between 9.9% to 24.5%, depending on the task.
(a) Entropy distributions across block draft heads
(b) Correlation between block efficiency and $h_{\max}$
Figure 3: (a) Entropy distributions across block draft heads on LAMBADA. The density plots illustrate the entropy distribution for each head in the model. (b) Correlation between block efficiency and $h_{\max}$, the head until which the average entropy in a task increases monotonically.
6.2 Confidence across multiple heads
Intuitively, predicting the identity of the $i^{\text{th}}$ future token becomes harder as $i$ increases. To better understand this phenomenon, we measure the confidence of the predictions by the entropy of the probability distribution. In 3(a), we plot the normalized histogram of entropy of each head on the LAMBADA task. From the normalized histogram, it is clear that the entropy increases as we move from first head to the last head, which agrees with our intuition that hardness of predictions increases as $i$ increases.
However, we observed that the entropy of heads does not increase monotonically for all tasks. Let ¯$\overline{\mathbb{H}}[i]$ be the average entropy of head $i$ on a particular corpus, and let $h_{\max}=\max\limits_{k}\{k:\forall i<k,\overline{\mathbb{H}}[i]\leq\overline{\mathbb{H}}[i+1]\}$, be the index of the largest head such that the average entropy of each head increases monotonically to that point. We observed a strong correlation between $h_{\max}$ and block efficiency (3(b)). Heads with lower entropy (indicating more confident predictions) intuitively contribute more to efficiency. Nonetheless, simply maximizing the number of low-entropy heads is not optimal, but rather incorporating progressively higher entropy heads, up to a certain point, can benefit decoding efficiency. A linear regression confirms this with an R-value of $\mathbf{0.77}$. This analysis suggests that BPD head entropy could be used as a proxy for block efficiency, and thus inference latency.
6.3 Oracle top-k block efficiency
Figure 4: An example of a top-5 sausage lattice generated on a NewsRoom example. Edge weights correspond to (rescored) logits. Edges at each time step are ordered in descending weight and green, bolded edges correspond to candidates matching the greedy decode over the next nine tokens: "… desktop computers with new Intel Corp processors that it …". The initial node in this graph is state 0 and the final node is 9.
Oracle efficiency  The concept of oracle block efficiency in BPD serves as a theoretical benchmark, illustrating the headroom available from improving the quality of the block draft. To compute oracle block efficiency, we consider the top-$k$ most probable tokens at each head, and form a “sausage” lattice from these. This data structure is a weighted directed graph, which succinctly represents all possible drafts (and their score under the BPD model) that could be formed from selecting one of $k$ tokens from each of the $h$ heads (Figure 4). In the automatic speech recognition and machine translation communities, it is known as a “confusion network” [36, 23].
Figure 5: Illustration of the output through oracle selection. For a given top $k$ of 3, if we can choose the oracle path successfully, the block efficiency can be improved from 1 to 5.
Given the top-$k$ lattice at each decoding step, we identify an oracle path that represents the path through the lattice that maximizes the length of the accepted prefix. This exercise, as shown in Figure 5, gives us insight into how much headroom exists in improving block drafts.
(a) LAMBADA
(b) SQuAD V1
(c) CNN/Daily
(d) SAMSUM
(e) MultiNews
(f) XSUM
Figure 6: Oracle block efficiency over the top-$k$ lattice as a function $k$. Each plot (a-f) represents a different task, demonstrating the relative improvement in block efficiency of the oracle draft with respect to the standard block draft as a function of the number of block draft heads used.
Potential headroom from oracle selection  Oracle drafting is not practical, but rather a reference point. Analyzing the gap between actual BPD performance and the oracle upper bound (Figure 6) helps us to understand the limitations of the original block drafts and potential areas for improvement. Additionally, exploring oracle efficiency as a function of the $k$ in the top-$k$ lattice, demonstrates how “close” the block draft was to producing a stronger draft.
7 Lattice rescoring for improved block efficiency
Having explored BPD’s prediction dynamics, we propose two drafting algorithms to improve block efficiency through rescoring of the top-$k$ lattice. This section presents techniques for rescoring the top-$k$ lattice along with empirical results.
Each of these algorithms is a modification of the block drafted in Stage 1 in Algorithm 1. Instead of using the most likely token at each head as the prediction, we construct the top-$k$ sausage lattice of likely drafts from each head, where the set of top-$k$ tokens is denoted as $S_{i}$ for head $i$. This approach allows any token within $S_{i}$ to be chosen for position $i$, yielding a total possible combinations of:
$|S_{1}|\times|S_{2}|\times\ldots|S_{h}|=k^{h}$
In this lattice, any path from the start to final state represents a viable draft. Two algorithms are proposed to select a small number of $h$-length drafts from this lattice, which are then passed to the verification step. The first algorithm employs neural autoregressive transformers (subsection 7.1), while the second utilizes n-gram language models (subsection 7.2).
[…]
Task Dataset Baseline Local rescoring Global rescoring Oracle (k=16)
BPD neural-61M BPD 4-gram BPD 16-best 0-gram BPD 16-best 4-gram BPD
[…]
7.1 Local rescoring via neural models
10:  : Blockwise parallel LM $\mathcal{M}_{\theta}^{h}$, top-$k$ indices selection function Top-$k(\cdot)$, rescoring model $\mathcal{M}_{\theta_{r}}$, interpolation weight $\alpha>0$.
21:  $z^{1}_{t+1},\dots,z^{h}_{t+h}\leftarrow\mathcal{M}_{\theta}^{h}(y_{t+1},\dots,y_{t+h}|\bar{x},y_{\leq t})$
32:  $S_{1},\dots,S_{h}\leftarrow\textsc{Top-}k(z^{1}_{t+1},\dots,z^{h}_{t+h})$
43:  /* Lattice Rescoring (subsection 7.1)*/
4:  for $j\leftarrow 2,\dots,h$ in parallel do
5:    $r_{t+j}\leftarrow\mathcal{M}_{\theta_{r}}(y_{t+j}|x,y_{<t+j})$
6:    $z^{j}_{t,j}[S_{j}]\leftarrow z^{j}_{t,j}[S_{j}]+\alpha\cdot r_{t+j}[S_{j}]$
7:  end for
Algorithm 2 Local rescoring via neural models
A simple approach uses a small neural rescorer, interpolating between the logits of the rescorer LM and vanilla block draft logits with an interpolation weight (algorithm 2). The rescored prediction is given by:
$z^{j}_{t,j}[S_{j}]\leftarrow z^{j}_{t,j}[S_{j}]+\alpha\cdot r_{t+j}[S_{j}]$
where $z^{j}_{t,j}$ represents the logit of the block draft at head $j$, and $r_{t+j}$ is the corresponding logit predicted by the small neural rescoring model, which is conditioned on the sequence $y_{\leq t},\cdots,\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1}$. The parameter $\alpha$ is the weight placed on the rescorer’s prediction. We experiment with decoder-only transformers having 32, 61, and 94 million (M) weight parameters (Appendix C). We use greedy rescoring when generating the neural draft.
7.2 Global n-gram rescoring
We also evaluate the quality of drafts generated by rescoring with an n-gram LM. Recall that blockwise parallel LMs can be used to compute a lattice representing $k^{h}$ possible sequences. We rescore all of these sequences with an n-gram  model, select the top $p$ sequences and pass them to the verification stage. When $p=1$, we refer to this as n-gram rescoring and when $p>1$, we refer to this as $p$-best n-gram BPD.
While global rescoring typically yields better results compared to local rescoring, rescoring $k^{h}$ sequences with a neural LM and selecting the most likely sequence would take time $O(k^{h})$, which is computationally prohibitive in most cases. Hence, we take advantage of n-gram models, which are unique in that we can select the most likely sequence in time $\text{poly}(k,h)$ using dynamic programming. We use the OpenFST library [1] to represent each n-gram model as a weighted finite state automaton and apply finite state composition with the top-$k$ lattice followed by extraction of the $p$ most likely draft sequences. Training details for the n-gram models are given in Appendix C.3.
(a) LAMBADA
(b) SQuAD V1
(c) CNN/Daily
(d) SAMSUM
(e) MultiNews
(f) XSUM
Figure 7: Block efficiency of $p$-best n-gram BPD methods as a function of the number of top $p$ sequences verified in parallel. The block efficiency of the methods is evaluated with the the same number of paths extracted from the top-$16$ lattice.
7.3 Empirical evaluation
Block efficiencyTable 4 and Figure 7 demonstrate the impact of lattice rescoring on block efficiency across various tasks. Autoregressive neural, n-gram LM, and $p$-best n-gram BPD rescoring all demonstrate improvements in block efficiency, although gains are task-dependent.
  • High initial block efficiency (LAMBADA, CNN/Daily): Both rescoring methods show little to no improvement, suggesting that vanilla BPD already produces high quality drafts.
  • Low initial block efficiency (SQuAD V1, SAMSUM, XSUM, NewsRoom): Both neural and n-gram rescoring lead to block efficiency gains, particularly with neural LMs achieving the best performance in some cases. This suggests that rescoring helps refine predictions and navigate the lattice more efficiently in these scenarios.
Table 5: Wins, ties, and losses of 61M parameter neural-rescored drafts and vanilla drafts. “Repair” corresponds to instances where the rescored draft eliminates repetition and “Regress” corresponds to instances where the rescored draft introduces repetition.
Dataset Ties Win Loss
Total % Repair % Regress Total % Repair % Regress
LAMBADA 631.5K 5804 27.95 0.05 9466 2.01 0.06
SQuAD V1 104.4K 1624 12.68 8.13 6325 2.53 12.28
CNN/Daily 965.0K 5928 23.20 0.67 17863 3.19 0.48
SAMSum 12.1K 2462 17.91 23.56 867 18.57 16.72
MultiNews 1.45M 294856 44.41 7.45 50209 22.21 5.37
XSUM 262.0K 36010 29.87 0.77 6826 4.19 10.99
NewsRoom 251.3K 79710 66.23 0.60 6492 2.85 7.39
Repairing repetitions  In subsection 6.1, we note that vanilla block drafts are prone to token-level repetition and that rescoring with a simple language model reduces the incidence of this. Although rescoring reduces repetition overall in drafts, is this driving improvements in block efficiency? To answer this, we compared the drafts generated by greedy rescoring with the 61M parameter neural rescorer against vanilla drafts. Time step instances were considered wins/ties/losses based on the accepted prefix length of the rescored draft vs. vanilla draft. Table 5 displays the win frequency across tasks along with the percentage of wins/losses attributed to introducing/eliminating repetition.
Note that in the tasks where rescoring improves block efficiency the most, NewsRoom and MultiNews, a high percentage of those repaired instances are driven by fixing erroneously repeated tokens. In fact, for MultiNews, 66.23% of block drafts are improved through repetition repair. We also evaluated the performance of rescoring with in-domain trained rescoring LMs, but found that they tended to perform no better than C4-trained LMs (Appendix D).
8 Conclusion
This paper presents a comprehensive analysis of BPD, highlighting its predictive dynamics and proposing methods to refine the generation of block drafts. Our study offers insights into BPD’s behavior, particularly the tendency for drafts to contain consecutive repetitions and its heads to exhibit varying confidence levels in predictions. We introduce a novel measure, oracle top-$k$ block efficiency, to explore potential improvements in block efficiency. Two algorithms are proposed for generating higher quality drafts: one for local rescoring with small neural models (i.e., neural BPD) and another for global rescoring with an n-gram LM and generating multiple drafts (i.e. $p$-best n-gram BPD). These algorithms leverage the strengths of both blockwise parallel LMs and small rescoring models to reduce average decoding latency, pushing the boundaries of efficient text generation with BPD. We believe that this paper lays the groundwork for future exploration in optimizing LM decoding speed.
Appendix A Broader Impact
Our research in advancing blockwise parallel decoding (BPD) for language models (LMs) paves the way for substantial improvements in language processing. This section outlines the broader implications of our work, focusing on key areas such as efficiency, scalability, and systematic impact.
  • Inference efficiency: We achieve further acceleration in inference with minimal increases in CPU overhead and memory bandwidth utilization. In particular, we show that our methods significantly boost the inference efficiency compared to the standard autoregressive inference.
  • Scalability: The introduction of selecting top $p$ sequences from top $k$ predictions demonstrates notable scalability, hinting at accelerated inference with growing model sizes. This scalability is further explored through diverse experiments involving various n-gram  and neural LM architectures.
  • Systematic impact: The advancement of BPD promises identical generation quality to traditional autoregressive methods, but with significantly improved efficiency in decoding times, which can be helpful in latency-critical applications.
Appendix B Limitation & Future Work
B.1 Limitation
Extending beyond greedy decoding
In this work, we focus on refining block drafts for greedy decoding, the decoding scheme considered in the original BPD paper [38]. Extending this methodology to non-greedy sampling strategies is an area for future work. Leviathan et al. [24] demonstrates the feasibility of applying speculative decoding strategies to accelerate (non-greedy) sampling. Their approach underscores the universality of the verification process across diverse model architectures, contingent on the availability of draft token-level predictions and probabilities — a requirement met by (rescored) block candidates.
Scaling up to larger models
We show that lattice rescoring can improve the block efficiency for a 1.5 billion parameter blockwise parallel LM. It remains an open question as to whether one would observe similar improvements in block efficiency with larger LMs.
Memory overhead during parallel verification
A notable challenge in our framework is the memory overhead encountered during parallel verification. The simultaneous processing of multiple token predictions, particularly for larger models and draft batches when coupled with lattice rescoring, can introduce high demands on memory. That said, this can be mitigated by the application of tree attention during verification [37].
Naive drafting heads and training recipes
The implementation of drafting heads within our BPD framework, while hewing close to the original BPD proposal, leaves room for architecture experimentation. The choice of architecture for the block drafter is crucial, and will likely affect the gains one observes from rescoring the block draft top-$k$ lattice. Future iterations could benefit from more sophisticated training methods, enhancing these heads’ ability to navigate complex linguistic contexts and improving overall predictive performance.
B.2 Future Work
The evolution of the BPD framework will be pivotal in addressing its current constraints and broadening its utility. Key areas for future research include:
  • The intrinsic compatibility of our lattice rescoring method with alternative sampling strategies presents fertile ground for future work. This delineation not only enriches the discourse within efficient language model inference but also sets a trajectory for subsequent empirical endeavors.
  • Scaling the blockwise parallel LM for compatibility with larger-scale LLMs, ensuring it remains effective and efficient as language models evolve.
  • Advancing training methodologies for drafting heads, to bolster their predictive accuracy and contribution to the BPD process, thereby optimizing the framework for more complex language modeling tasks.
  • Using the sequential entropy ordering of heads (3(b)) as a possible halting condition during block draft head training, or to inform how a rescoring LM should be interpolated with the block lattice weights.
Appendix C Experiment Details
C.1 Training objective for blockwise parallel LMs
We minimized the following loss function to train blockwise parallel LMs:
$\mathcal{L}_{BPD}=\sum_{h=1}^{H}\lambda_{h}\mathcal{L}_{h}$
where $H$ is the number of heads, $\lambda_{h}$ is a non-negative scalar that weights the loss from head $h$, and $\mathcal{L}_{h}$ denotes the loss for each individual head:
$\mathcal{L}_{h}=-\sum_{x_{1\ldots i},y_{i+h}}\log p(y_{i+h}|x_{1\ldots i})$
where $x_{1\ldots i}$ is the token sequence up to position $i$, $y_{i+h}$ is the ground truth token at position $i+h$, and $p(y_{i+h}|x_{1\ldots i})$ is the probability of observing token $y_{i+h}$ given the sequence $x_{1\ldots i}$ under the blockwise parallel LM. We trained all models in this work with $\lambda_{h}=1$. We leave tuning these hyperparameters, improving the block efficiency and quality of the blockwise parallel LM, as future work.
C.2 Neural Model Details
Table 6: Architecture hyperparameters for each of the transformer-based neural language models.
Type Model # Layers Embedding Dim Hidden Dim
Blockwise Parallel Decoder 1.5B 18 1,536 12,288
Autoregressive Decoder 32M 2 384 1,536
61M 12 384 1,536
94M 6 768 3,072
Each neural rescoring LM is a decoder-only transformer with learned absolute positional embeddings and twelve self-attention heads at each layer. The key architecture hyperparameters are given in Table 6. Aside from scale, the only difference between the blockwise parallel LM and neural rescoring models is the addition of the feedforward neural networks and eight additional block prediction heads. Note that the number of parameters for each of these models also includes the embedding table.
Each model was pretrained on the English C4 corpus for 200K iterations with a batch size of $2^{20}\approx 1M$ tokens per batch. Dropout was not applied. For the blockwise parallel LM, all heads were trained jointly. The pretraining for the blockwise parallel LMs took about 47 hours on 128 TPUv3 units.
For downstream tasks, models were finetuned for a maximum 100K iterations with a batch size of two examples with maximum sequence length of 2048. Maximum learning rate was fixed to $10^{-4}$ for all runs, with a cosine learning rate schedule. Checkpoints were selected based on heldout set model performance. Interpolation weight for all rescoring models was tuned for block efficiency on 100 randomly selected examples from the evaluation set for each task, and performance was reported on the remainder of the evaluation set.
C.3 n-gram Details
All n-gram  LMs in this work are Katz backoff n-gram LMs [19] fit on the train split of the GPT3 subword-tokenized English C4 corpus with n-gram order $\in\{2,4\}$. We apply entropy pruning [39] to reduce model size to a maximum of 100 million n-grams per model, and ensure that each trigram is observed at least twice and each 4-gram is observed at least four times. Preprocessing of the text is identical to that used to train neural LMs.
C.4 Datasets
  • LAMBADA (LAnguage Modeling Broadened to Account for Discourse Aspects): A collection of narrative passages designed to test the understanding of long-range dependencies in language models, where the task involves predicting the last word of a passage based on the full context [29].
  • SQuAD V1 (Stanford Question Answering Dataset): A reading comprehension dataset that features questions based on Wikipedia articles, with answers located within the text [34].
  • CNN/DailyMail: This dataset includes news articles paired with human-written summaries, mainly used to evaluate the summarization capabilities of language models, particularly in abstractive summarization [15].
  • SAMSum (Semi-Automatic Machine Summarization): Focuses on abstractive summarization using news articles and machine-generated summaries, testing models’ abilities to refine and improve existing summaries [12].
  • MultiNews: Comprises news articles from diverse sources for abstractive summarization tasks, evaluating models on handling different writing styles and topics [11].
  • XSUM: Contains scientific documents and summaries, challenging language models to process complex scientific information and language [28].
  • NewsRoom: A dataset of news articles aimed at assessing the factual accuracy and information extraction capabilities of models in generating summaries [13].
All datasets were tokenized using the 50,257 GPT3 subword vocabulary [3].
Templates
We used the following prompts during model finetuning and inference.
  • SQuAD: "question: [question] context: [context]"
  • CNN/DailyMail: "summarize: [text]"
[…]
  • MultiNews: "Write a summary based on this article: [text]"
  • XSUM: "Summarize: [text]"
  • NewsRoom: "Please write a short summary for the following article: [title] [text]"
Appendix D Rescoring with in-domain language models
Table 7: Block efficiency from rescoring with in-domain trained rescoring models for 2-gram and 61M parameter neural rescorer.
Dataset 2-gram neural-61M
C4 In-domain C4 In-domain
SQuAD V1 2.09 2.04 2.10 2.06
CNN/Daily 1.73 1.73 1.73 1.72
SAMSUM 1.31 1.22 1.39 1.24
MultiNews 1.13 1.14 1.25 1.16
XSUM 1.17 1.18 1.23 1.14
NewsRoom 1.20 1.22 1.29 1.11
We found that in-domain rescorers performed no better than rescorers only trained on C4. We suspect this is due to a lack of sufficient finetuning data and that the main benefit of rescoring comes from discouraging unnatural artifacts such as repetition from the original BPD draft. Table 7 shows block efficiency after rescoring using in-domain models for all tasks besides language modeling.
Neural rescorers were finetuned from C4-pretrained checkpoints. n-gram  models were trained from scratch, and unseen vocabulary was added as unigram arcs with trivial weight (negative log probability of 1000.0). This was done to ensure that all paths through the lattice were assigned non-zero probability by the n-gram  model. In previous experiments, we also tried interpolating the in-domain n-gram  model with a unigram model trained on C4, and observed similar performance as simply adding unseen unigrams.
Appendix E Interpolation weights tuned per task
We tuned the interpolation weight, $\alpha$ for the 94M parameter neural and 4-gram LM rescorers, and then used this weight to rescore with all other models of that same class. 100 examples from each task’s heldout set were set aside for tuning, to maximize block efficiency. The remainder of examples were used for evaluation. We swept over $\alpha\in\{0.1,0.5,0.75,0.9,1.0,1.1,1.5,2.0,5.0,10.0\}$.
Note that for tasks where lattice rescoring was unhelpful, the interpolation weight, $\alpha$ is tuned to place much higher weight on the block draft logits (Table 8). This is a signal that the rescorer does not provide additional information over the original block draft heads.
Table 8: Tuned interpolation weight per task for neural and n-gram rescoring.
Dataset Neural n-gram
LAMBADA 0.1 0.1
SQuAD V1 1.0 0.75
SAMSum 5.0 1.5
CNN/Daily 0.1 0.1
MultiNews 5.0 2.0
XSUM 1.5 1.1
NewsRoom 5.0 2.0
Appendix F Local rescoring impact on block efficiency
[…]
Task Dataset Baseline Global rescoring Local rescoring Oracle (k=2) Oracle (k=16)
BPD 2-gram BPD 3-gram BPD 4-gram BPD neural-32M BPD neural-61M BPD neural-94M BPD
[…]
Table 9 reveals the impact of different rescoring methods on the block efficiency of the block lattice, offering valuable insights into their effectiveness across diverse tasks and models, supporting the investigations in subsection 7.1.
  • Limited improvement for high baselines: For tasks with already high initial block efficiency (LAMBADA, CNN/DailyMail), rescoring offers minimal or even negative changes in block efficiency compared to the baseline BPD system. This suggests that for tasks where standard BPD already achieves significant speed improvements, there is limited room for further gains through rescoring.
  • Efficacy for poor baselines: In tasks with lower initial block efficiency (SQuAD V1, XSUM, NewsRoom), rescoring using both n-gram and neural language models results in increased block efficiency. Notably, neural rescoring with larger models (61M and 94M parameters) achieves the highest efficiency gains in these tasks, reaching up to 19.44% improvement in NewsRoom. These results highlight the potential of rescoring to refine predictions and enhance efficiency for models exhibiting calibration issues.
  • Task-specific effectiveness: The level of improvement from rescoring varies across different summarization tasks (MultiNews, XSUM, NewsRoom). While all show positive responses, NewsRoom exhibits the largest gains, suggesting that the effectiveness of rescoring is task-dependent.
  • Comparison with oracle efficiency: The ‘Oracle’ columns present the upper bound achievable if only the most likely token at each step is chosen with perfect hindsight (k=2 and k=16). While significant gaps remain between current results and the oracle, the observed improvements from rescoring demonstrate progress towards closing this efficiency gap.
Overall, these findings suggest that local rescoring methods can be a valuable tool for enhancing BPD efficiency, particularly for models with less calibrated predictions. Further exploration of advanced rescoring strategies, especially in conjunction with larger neural language models, holds promise for achieving even closer-to-oracle efficiency levels.
Appendix G Ablation on the number of heads in the blockwise parallel LM
Table 10 summarizes the block efficiency for different head configurations across various language tasks with the same settings discussed in Table 1.
  • General trend: Both performance and block efficiency tend to increase with the number of heads, up to a point. This suggests that using more heads allows the model to capture richer contextual information and make more accurate predictions.
  • Efficiency trade-off: While increasing heads generally improves block efficiency, it also increases the memory for verification stages. Therefore, the optimal number of heads depends on the balance between desired block efficiency and available resources.
Table 10: Test performance per task. Test performance of each finetuned model and block efficiency are shown as a function of heads ($h\in{3,6,9}$). Tasks inclue Language Modeling (LM), extractive Question Answering (QA), and both Long and Short Summarization (L-Sum & S-Sum). The metric for LM is perplexity, for QA is exact match, and for all the remaining (summarization) tasks, the metric is ROUGE-L.
Task Dataset Performance # of Heads ($h$)
3 6 9
LM LAMBADA 7.88 1.79 2.84 3.12
QA SQuAD V1 57.60 1.53 2.03 2.08
S-SUM CNN/Daily 39.85 1.60 1.71 1.74
SAMSUM 37.66 1.18 1.25 1.27
L-SUM MultiNews 23.08 1.08 1.08 1.10
XSUM 52.15 1.11 1.12 1.13
NewsRoom 39.85 1.07 1.08 1.08
Appendix H Practical efficiency of rescoring block drafts
To enhance our understanding of block rescoring within the realm of contemporary deep learning hardware environments, we present an in-depth examination focused on TPU/GPU utilization and the overhead incurred by n-gram  rescoring. This analysis is divided into two parts: (1) an analysis of block rescoring through the lens of TPU/GPU utilization, and (2) empirical benchmarks of n-gram  lattice rescoring. The major takeaways are as follows.
Memory bandwidth (HBM $\Leftrightarrow$ SRAM)
A critical factor in the performance of deep learning applications is the efficient management of memory bandwidth between High Bandwidth Memory (HBM) and Static Random Access Memory (SRAM) [8]. Increasing the block efficiency via the block lattice rescoring reduces the average per token parameter and key-value cache I/O that needs to be communicated from HBM to SRAM.
Overhead in n-gram  rescoring
n-gram  rescoring is actually quite efficient. For the size of lattices we consider in this work, moving the lattice from HBM to DRAM, performing n-best n-gram  rescoring, and moving the n-best paths back to HBM requires no more than 2 ms per lattice.
H.1 Hardware utilization
We compare our approach against traditional Autoregressive LMs across several metrics (Table 11):
Table 11: Comparative analysis of per decoded token efficiency metrics across block rescoring methods and the standard autoregressive LM (batch size=1). This table shows the average block efficiency, parameter I/O, key-value (KV) cache I/O at varying sequence lengths, and FLOPS—evaluated on a per-token basis with batch size 1.
Component Autoregressive Base BPD 4-gram BPD Neural-61M BPD 16-best 0-gram BPD 16-best 4-gram BPD
Avg. Block Efficiency 1.000 1.646 1.657 1.724 1.717 1.797
Parameter I/O (GB) 3.000 1.823 1.811 1.811 1.747 1.669
KV Cache I/O (GB) - Seq_len 128 0.113 0.074 0.073 0.076 0.140 0.134
KV Cache I/O (GB) - Seq_len 512 0.453 0.280 0.278 0.290 0.338 0.323
KV Cache I/O (GB) - Seq_len 1024 0.906 0.555 0.552 0.574 0.602 0.575
KV Cache I/O (GB) - Seq_len 2048 1.812 1.106 1.098 1.144 1.129 1.079
FLOPS (T) 0.931 0.57 0.567 0.635 0.621 0.593
Memory bandwidth and compute efficiency
The block rescoring variants achieve significant reductions in Parameter I/O and KV Cache I/O compared to autoregressive decoding, suggesting BPD methods’ ability to reducing inference times by mitigating the primary latency bottleneck—memory bandwidth. Advances in TPU/GPU architecture ensure that an increase in FLOPS per token has a minimal effect on latency, confirming our strategy’s capacity to navigate the complexities of memory bandwidth efficiently.
Comparative latency impact
A consistent decrease in memory bandwidth utilization across blockwise parallel LMs, including those leveraging LM rescoring and parallel processing strategies, illustrates our approach’s contribution to accelerating inference speed. This underscores the practicality and applicability of our enhancements in promoting more efficient language model inference within state-of-the-art computational frameworks.
H.2 Overhead of n-gram rescoring
While the majority of computational efforts in block rescoring are dedicated to TPU/GPU utilization, the implementation of n-gram rescoring introduces additional overheads. These are primarily attributed to CPU computations and the data transfer between the CPU and HBM. This section provides a comprehensive examination of these overheads, drawing on benchmarks from rescoring experiments with a 4-gram C4 LM.
Benchmarks for 4-gram C4 LM rescoring
We conducted benchmarks on rescoring lattices with a 4-gram C4 LM of $\approx$100M n-grams. The average latency observed across 10 runs for different numbers of the shortest paths is summarized in Table 12.
Table 12: Average latency for N-best rescoring an 8-time step lattice with 16 arcs per time step. N, the number of shortest paths, is varied from 1 to 16.
# Shortest Paths N-best Rescoring Latency (ms)
1 1.630
2 1.751
4 1.878
8 1.871
16 1.983
Notably, rescoring with a large 4-gram LM averages less than 2 milliseconds for extracting up to 16 globally-best paths, despite the lattice containing approximately 4.29 billion possible paths. In our initial experiments, increasing the size of the n-gram  LM had little effect on n-best rescoring latency, indicating that improvements to rescoring LM quality will incur little additional latency, provided that the rescoring LM fits within DRAM.
Latency is predominantly influenced by lattice size, particularly the number of top-k tokens per time step and the number of time steps, as depicted in Table 13.
Table 13: 1-best rescoring latency by the 4-gram C4 LM for varying lattice sizes.
Number of time steps Top-k per time step 1-best rescoring latency (ms)
4 2 1.038
4 4 1.050
4 8 1.130
4 16 1.237
8 2 1.061
8 4 1.144
8 8 1.234
8 16 1.630
16 2 1.102
16 4 1.206
16 8 1.558
16 16 2.174
The benchmarks highlight the fact that the additional overhead introduced by n-gram  rescoring, though present, should not significantly impact overall latency.